Find the quadratic equation of the roots:
5+2i and 5-2i

Answer by mananth(16946)

5+2i and 5-2i
let alpha beta be the roots of the equation
sum of roots= 5+2i + 5-2i= 10
product of roots = (5+2i)(5-2i)
=25-*(2i)^2
=25 -4i^2
=25 -4*-1
=25+4
=29
the equation is x^2-10x+29=0